OpenStudy (anonymous):

Suppose that f(x)=4x+7 and g(x)=2x. What is a rule for g(f(x))? What is a rule for f(g(x))

OpenStudy (anonymous):

g(f(x))=g(4x+7)=2(4x+7)=8x+14 f(g(x))=f(2x)=4(2x)+7=8x+7